OpenStudy (anonymous):

Find the surface area of a prism with a square base that is 5 cm by 5 cm and is 10 cm tall. A. 250 cm2 B. 785 cm2 C. 250 cm3 D. 100 cm2

OpenStudy (anonymous):

@k_lynn i am also stuck on this one, im not sure what they mean by prism with a square base so i dont know which formula to use.

OpenStudy (k_lynn):

Do you have a picture you can show me?

OpenStudy (anonymous):

@k_lynn it does not come with a picture, thats why i am confused.

OpenStudy (lsdxmdmaddicthc):

Surface Area = (Area of Base * 2) + (Lateral Area)

OpenStudy (lsdxmdmaddicthc):

Area of Base would be the Area of a Square. Area of a Square = S*S (5*5)

OpenStudy (automaticloveletter):

SA=2*a^2+4*a*h A= Base Edge (5cm) H= Height (10cm) SA=250

OpenStudy (k_lynn):

http://prntscr.com/6z7735 I'm assuming since the question says prism, it looks like this.

OpenStudy (lsdxmdmaddicthc):

Lateral Area = The height of the prism * the perimeter of the base

OpenStudy (anonymous):

So, Surface area= (25 * 2) + 103.08 SA= 50 +103.08 SA= 153.08 this is what i got but that isn't any of the choices:/

OpenStudy (lsdxmdmaddicthc):

Height of the prism = 10 Perimeter of the Base = 5 + 5 + 5 + 5

OpenStudy (k_lynn):

I think what we're all trying to say is that when you find the surface area, you find the area of all the sides, and then add them together.

OpenStudy (lsdxmdmaddicthc):

Tayler, follow me :)

OpenStudy (lsdxmdmaddicthc):

Okay so the surface area is the sum of the areas of both bases, plus the Lateral Area

OpenStudy (lsdxmdmaddicthc):

SA = (BA*2) + Lateral Area Base Area = s*s BA = 5*5 = 25 Lateral Area = Perimeter of base * The height of the prism Perimeter of base = 5 + 5 + 5 + 5 Height of prism = 10

OpenStudy (lsdxmdmaddicthc):

I think you can take it from here ;)

OpenStudy (lsdxmdmaddicthc):

Sometimes it can help ALOT to draw it out on scrap paper

OpenStudy (anonymous):

@lsDxMdmaddicThc thanks so much!:)

OpenStudy (lsdxmdmaddicthc):

Thank you for learning and not just copying down an answer!

OpenStudy (lsdxmdmaddicthc):

Glad I could help you! :)